OpenStudy (anonymous):

What is the equation of the line, in point-slope form, that passes through the points (-3, -1) and (-6, 8)? y + 8 = -3(x - 6) y - 8 = -3(x + 6) y + 8 = 3(x - 6) y - 8 = 3(x + 6)

TheSmartOne (thesmartone):

The slope of two points: \(\sf (x_1,y _1)\) and\(\sf (x_2,y_2)\) is: \(\sf\Large Slope=\frac{y_2-y_1}{x_2-x_1}\)

TheSmartOne (thesmartone):

And why is the point-slope form?

OpenStudy (anonymous):

idk

TheSmartOne (thesmartone):

Ok, then google "point-slope form" and tell me what it is.

OpenStudy (anonymous):

Point slope form is also the quickest method for finding the equation of line given two points

TheSmartOne (thesmartone):

what equation does it show you that you get for point-slope form?

OpenStudy (anonymous):

idk im bad at math

TheSmartOne (thesmartone):

Not an excuse... Read the first 2 paragraphs: http://www.purplemath.com/modules/strtlneq2.htm

TheSmartOne (thesmartone):

Point slope form is: \(\sf y – y_1 = m(x – x_1)\)

OpenStudy (anonymous):

ok?

TheSmartOne (thesmartone):

calculate the slope first...

OpenStudy (anonymous):

idk how

TheSmartOne (thesmartone):

Read my first reply on your post, it shows you how to.

OpenStudy (anonymous):

i got 2.3 for theslope

TheSmartOne (thesmartone):

How? Where is the work?

TheSmartOne (thesmartone):

*facepalm* You used a calculator, yet you plugged in the wrong point...

OpenStudy (anonymous):

my name explains me man

TheSmartOne (thesmartone):

this will give you the correct slope... http://www.calculator.net/slope-calculator.html?type=1&x11=-3&y11=-1&x12=-6&y12=8&x=92&y=4

OpenStudy (anonymous):

now what

TheSmartOne (thesmartone):

You have the slope, so now you can plug it into: \(\sf y – y_1 = m(x – x_1)\) m=slope \(\sf (x_1,y_1)\) is a point what is the slope again?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

can u just tell me answer

TheSmartOne (thesmartone):

\(\sf (-3, -1)\) and \(\sf (-6,8)\) are both in the form of \(\sf (x_1,y_1)\) substitute any point into the equation.